The following fractions represent just three different numbers. Separate them into three groups of equivalent fractions, by changing each one to its simplest form.
15/75

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

15 / 75 = (1 × 15) / (5 × 15)
=1 / 5

​Thanks